OpenStudy (anonymous):

@dumbcow In a lottery, there are 2 000 000 tickets to be sold. The prizes are as follows: 1x$1000000 5x$50000 10x$1000 50x$50 What should the lottery operators charge per ticket in order to make a 40% profit?

OpenStudy (anonymous):

a

OpenStudy (dumbcow):

not quite sure what the 40% is referring to profit =sales - prizes \[\text{profit} = 2,000,000 x - 1,262,500\]

OpenStudy (anonymous):

Yea thats what i dont get

OpenStudy (anonymous):

I think its revenue

OpenStudy (dumbcow):

hmm ok lets just base it on the prize money then if they give the tickets away they lose 100% of their money so set profit as 40% of total prize money, then solve for x

OpenStudy (anonymous):

the answer in the book is $1.05

OpenStudy (anonymous):

I keep getting 88 cents

OpenStudy (anonymous):

This is what someone else did but i dont under stand why 60% of revenue is the prize money 40% of what? Cost, or revenue? prize money = 1,000,000×1 + 50,000×5 + 10,000×10 + 50×50 = 1,252,500 prize money = 60% of revenue revenue = 1,252,500/60% = 2,104,167 2,104,167 dollars/2,000,000 tickets = 1.052083 dollars/ticket https://ca.answers.yahoo.com/question/index?qid=20131129082240AAXx5sD

OpenStudy (anonymous):

can you explain if you get how they got it?

OpenStudy (dumbcow):

oh ok now i get it... what they mean is that 40% profit refers to money they keep after giving out prize money so they keep 40% of revenue and give out 60%

OpenStudy (dumbcow):

\[(.6)(2,000,000 x) = 1,262,500\]

OpenStudy (anonymous):

oh okay, i c

OpenStudy (anonymous):

Thanks for explaining :)

OpenStudy (dumbcow):

no problem